In a household during a particular week three 2kW fires are used on average 25h each and eight 100W light bulbs are used on average 35h each.Determine the cost of electricity for the week if 1 unit of electricity costs 15p.
1
Expert's answer
2021-09-23T08:30:12-0400

First, we have to calculate the energy consumed in total and it has to be expressed in kWh, thus the consumption of the light bulb will be 100 W=0.100 kW, then we find:


"\text{Power consumed}=(3\,fires)(2\frac{kW}{fire})(25\,h)+(8\,bulbs)(0.100\frac{kW}{bulb})(35\,h)
\\ \text{Power consumed}={(6\times25+7\times4)}\,kWh={(150+28)}\,kWh=178\,kWh"


After that, we can calculate the total cost by just multiplying two factors:


"\\ \text{Cost of electricity}=(178\,\cancel{kWh})\Large{(\frac{15p}{\cancel{kWh}})}=2670\,p"


In conclusion, the total cost of electricity will be 2670 p.
Reference:

  • Sears, F. W., & Zemansky, M. W. (1973). University physics.